### v2

# Dynamical degrees of automorphisms of K3 surfaces with Picard number 22

Abstract
We show that there exists an automorphism of a projective K3 surface with Picard number $2$ such that the trace of its action on the Picard lattice is $3$. Together with a result of K. Hashimoto, J. Keum and K. Lee, we determine the set of dynamical degrees of automorphisms of projective K3 surfaces with Picard number $2$.
footnotetext: MSC(2020): 14J28, 14J50, 11H56.
1 Introduction
Let $S$ be a compact Kähler surface. The intersection form makes $H^{2}(S,\mathbb{Z})_{f}$ a unimodular lattice, where $H^{2}(S,\mathbb{Z})_{f}\coloneqq H^{2}(S,\mathbb{Z})/\textrm{torsion}$. Let $\phi$ be an automorphism of $S$. Then, the induced homomorphism $\phi^{*}:H^{2}(S,\mathbb{Z})_{f}\to H^{2}(S,\mathbb{Z})_{f}$ is an isometry of the lattice $H^{2}(S,\mathbb{Z})_{f}$. Let $\lambda(\phi)$ denote the spectral radius of $\phi^{*}:H^{2}(S,\mathbb{C})\to H^{2}(S,\mathbb{C})$:
$\lambda(\phi)\coloneqq\max\{|\mu|\mid\text{$\mu\in\mathbb{C}$ is an eigenvalue of $\phi^{*}:H^{2}(S,\mathbb{C})\to H^{2}(S,\mathbb{C})$}\}.$
By the Gromov–Yomdin theorem, the topological entropy of $\phi$ is given by $\log\lambda(\phi)$. We call the value $\lambda(\phi)$ the (first) dynamical degree of $\phi$ (a general definition of dynamical degrees can be found in [4]). According to S. Cantat [3], if a compact complex surface admits an automorphism $\phi$ with $\lambda(\phi)>1$, then the surface is bimeromorphic to a torus, a K3 surface, an Enriques surface, or a rational surface.
A K3 surface is a compact complex surface $S$ such that it has a nowhere vanishing holomorphic $2$-form $\omega_{S}$ and its irregularity ($\dim H^{1}(S,\mathcal{O}_{S})$) is zero. Every K3 surface is Kähler [14]. Let $S$ be a K3 surface. The submodule $P_{S}\coloneqq H^{2}(S,\mathbb{Z})\cap H^{1,1}(S)$ of $H^{2}(S,\mathbb{Z})$ is called the Picard lattice or the Néron-Severi lattice. The Picard number $\rho_{S}$ is the rank of $P_{S}$. The second cohomology group $H^{2}(S,\mathbb{Z})$ is a free $\mathbb{Z}$-module of rank $22$ and, equipped with the intersection form, forms an even unimodular lattice of signature $(3,19)$. We refer to an even unimodular lattice of signature $(3,19)$ as a K3 lattice. It is known that such a lattice is unique up to isomorphism.
Let $\phi$ be an automorphism of $S$. It is known that $\lambda(\phi)$ is equal to either $1$ or a Salem number, that is, a real algebraic integer $\beta>1$ such that $\beta^{-1}$ is a conjugate of $\beta$ and all of its conjugates other than $\beta$ and $\beta^{-1}$ have absolute value $1$. The statement also holds for an automorphism of a compact Kähler surface, see [8, Theorem 3.2] and [11, Theorem 2.5]. A natural question is to determine which Salem numbers can appear as dynamical degrees of automorphisms of K3 surfaces. In this article, we determine the set
[…]
Assume that $S$ is a projective K3 surface. Then $\lambda(\phi)$ is given by the spectral radius of the restriction $\phi^{*}|_{P_{S}}$. Hence, the degree (as an algebraic number) of $\lambda(\phi)$ is at most $\rho_{S}$. Suppose further that $\rho_{S}=2$ and $\lambda(\phi)>1$. Then $\lambda(\phi)$ must be a Salem number of degree $2$. Its minimal polynomial is $X^{2}-(\lambda(\phi)+\lambda(\phi)^{-1})X+1$, and in particular, $\lambda(\phi)+\lambda(\phi)^{-1}\in\mathbb{Z}_{\geq 3}$. So, we consider the set
[…]
rather than $\mathcal{D}$. Note that $\mathcal{D}=\{(\tau+\sqrt{\tau^{2}-4})/2\mid\tau\in\mathcal{T}\}$.
In their paper [6], K. Hashimoto, J. Keum and K. Lee study the cases $\phi^{*}\omega_{S}=\omega_{S}$ and $\phi^{*}\omega_{S}=-\omega_{S}$, and prove the following theorem.
Theorem 1.1 ([6, Main Theorem]).
Let $\tau\geq 3$ and $\epsilon\in\{1,-1\}$. Put $A_{1}=\mathbb{Z}_{\geq 4}$ and $A_{-1}=\mathbb{Z}_{\geq 4}\setminus\{5,7,13,17\}$. The following are equivalent:
  1. (i)
    There exists an automorphism $\phi$ of a projective K3 surface $S$ with Picard number $2$ such that $\lambda(\phi)+\lambda(\phi)^{-1}=\tau$ and $\phi^{*}\omega_{S}=\epsilon\omega_{S}$.
  2. (ii)
    There exists $\alpha\in A_{\epsilon}$ such that $\tau=\alpha^{2}-2\epsilon$.
It follows from this theorem that
$\{2\}\cup\{\alpha^{2}-2\mid\alpha\in\mathbb{Z}_{\geq 4}\}\cup\{\alpha^{2}+2\mid\alpha\in\mathbb{Z}_{\geq 4}\setminus\{5,7,13,17\}\}\subset\mathcal{T},$
where $2\in\mathcal{T}$ since $\lambda(\mathrm{id})+\lambda(\mathrm{id})^{-1}=2$. This article fills the gap left by this inclusion.
Theorem A.
There exists a projective K3 surface $S$ with Picard number $2$ and an automorphism $\phi$ of $S$ such that the characteristic polynomial of the induced homomorphism $\phi^{*}:H^{2}(S,\mathbb{C})\to H^{2}(S,\mathbb{C})$ is $(X^{2}-3X+1)\cdot\Phi_{50}(X)$, where $\Phi_{50}(X)$ is the $50$th cyclotomic polynomial.
To prove this theorem, we will construct an isometry of a K3 lattice with characteristic polynomial $(X^{2}-3X+1)\cdot\Phi_{50}(X)$ by using methods in [9, 10]. Then, it leads to the existence of the desired automorphism of a K3 surface by the Torelli theorem and surjectivity of the period mapping.
Let $\phi$ be the automorphism in Theorem A. Then $\lambda(\phi)+\lambda(\phi)^{-1}=3$, which implies that $3=1^{2}+2\in\mathcal{T}$. Furthermore, we get $7=3^{2}-2\in\mathcal{T}$ because
$\lambda(\phi^{2})+\lambda(\phi^{2})^{-1}=\lambda(\phi)^{2}+\lambda(\phi)^{-2}=(\lambda(\phi)+\lambda(\phi)^{-1})^{2}-2=7.$
Hence,
$\{2\}\cup\{\alpha^{2}-2\mid\alpha\in\mathbb{Z}_{\geq 3}\}\cup\{\alpha^{2}+2\mid\alpha\in\mathbb{Z}_{\geq 1}\setminus\{2,3,5,7,13,17\}\}\subset\mathcal{T}.$ (2)
Moreover, some further discussions combined with Theorem 1.1 show that the inclusion is, in fact, an equality.
Theorem B.
We have
$\mathcal{T}=\{2\}\cup\{\alpha^{2}-2\mid\alpha\in\mathbb{Z}_{\geq 3}\}\cup\{\alpha^{2}+2\mid\alpha\in\mathbb{Z}_{\geq 1}\setminus\{2,3,5,7,13,17\}\}.$
Remark 1.2.
P. Reschke [12, 13] shows that every Salem number of degree $2$ is realized as the dynamical degree of an automorphism of a complex $2$-torus (in this case, the torus must be projective). Let $\psi$ be an automorphism of a $2$-torus AA whose dynamical degree is a given Salem number $\beta$ of degree $2$. Then $\psi$ gives rise to an automorphism $\phi$ of the Kummer surface obtained from $A$, and $\phi$ also has dynamical degree $\beta$, see [8, §4]. Hence, every Salem number of degree $2$ is realized as the dynamical degree of an automorphism of a projective K3 surface. However, we remark that the Picard number of a Kummer surface is greater than or equal to $16$.
Remark 1.3.
Let $S$ be a non-projective K3 surface and $\phi$ an automorphism of $S$ with $\lambda(\phi)>1$. Then $\lambda(\phi)$ is a Salem number of degree at least $4$. Furthermore, if $d$ denotes the degree of $\lambda(\phi)$, then $\rho_{S}=22-d$.
It is shown that every Salem number of degree $d$ is realizable as the dynamical degree of an automorphism of a non-projective K3 surface for $d=4,6,8,12,14,16$ by E. Bayer-Fluckiger [1], and for $d=20$ by the author [15] (a simple characterization in the case $d=22$ is given in [2]). In particular, as a non-projective analog of Theorem B, we have
[…]
2 Lattices
This section summarizes some known results on lattices. For more details on gluing and twist, we refer to [9].
Lattices and isometries
A lattice is a finitely generated free $\mathbb{Z}$-module $L$ equipped with an inner product, i.e., a nondegenerate symmetric bilinear form $b:L\times L\to\mathbb{Z}$. Let $L=(L,b)$ be a lattice. Its dual is defined to be
$L^{\vee}\coloneqq\{y\in L\otimes\mathbb{Q}\mid b(y,x)\in\mathbb{Z}\text{ for all $x\in L$ }\},$
where $b$ is extended on $L\otimes\mathbb{Q}$ linearly. Since $b$ is $\mathbb{Z}$-valued, we have $L\subset L^{\vee}$. The lattice $L$ is unimodular if $L=L^{\vee}$. We say that $L$ is even if $b(x,x)\in 2\mathbb{Z}$ for all $x\in L$. The signature of $L$ is the signature of the inner product space over $\mathbb{R}$ obtained by extending scalars to $\mathbb{R}$.
Let $e_{1},\ldots,e_{d}$ be a basis of $L$. The $d\times d$ matrix $G\coloneqq(b(e_{i},e_{j}))_{ij}\in M_{d}(\mathbb{Z})$ is called the Gram matrix of $(L,b)$ with respect to $e_{1},\ldots,e_{d}$. Its determinant $\det G$ does not depend on the choice of the basis. This integer is called the determinant of $L$ and denoted by $\det L$ or $\det b$. It is clear that $L$ is unimodular if and only if $|\det G|=1$, and that $L$ is even if and only if all diagonal entries of $G$ are even.
Let $(L_{1},b_{1})$ and $(L_{2},b_{2})$ be two lattices. An isometry $t:L_{1}\to L_{2}$ is a homomorphism of $\mathbb{Z}$-modules satisfying $b_{2}(t(x),t(y))=b_{1}(x,y)$ for all $x,y\in L_{1}$. When $L_{1}=L_{2}$, we say that $t$ is an isometry of $L_{1}$.
Theorem 2.1.
Let $L$ be an even unimodular lattice of rank $2n$, and let $F(X)\in\mathbb{Z}[X]$ be the characteristic polynomial of an isometry of $L$. Then $|F(1)|$, |F(1)||F(-1)|, and $(-1)^{n}F(1)F(-1)$ are squares.
Proof.
See [5, Theorem 6.1] or [1, Theorem 2.1]. ∎
Glue groups
Let $L=(L,b)$ be a lattice. The quotient $L^{\vee}/L$ is called the glue group or the discriminant group of $L$, and denoted by $G(L)$. It is a finite abelian group of order $|\det L|$. In fact, if $d_{1},\ldots,d_{r}\in\mathbb{Z}_{\geq 1}$ are the elementary divisors of its Gram matrix then $G(L)\cong(\mathbb{Z}/d_{1}\mathbb{Z})\oplus\cdots\oplus(\mathbb{Z}/d_{r}\mathbb{Z})$. For a prime $p$, let $G(L)_{p}$ denote the Sylow $p$-subgroup of $G(L)$. Then $G(L)$ decomposes as $G(L)=\bigoplus_{p}G(L)_{p}$ where $p$ ranges over the primes dividing $\det L$. We remark that if $p.G(L)_{p}=0$ then $G(L)_{p}$ is isomorphic to the direct sum of finitely many copies of $\mathbb{Z}/p\mathbb{Z}$ and can be regarded as an $\mathbb{F}_{p}$-vector space.
For $x\in L^{\vee}$, we write $\bar{x}=x+L\in G(L)$. The glue group admits a nondegenerate torsion form $\bar{b}:G(L)\times G(L)\to\mathbb{Q}/\mathbb{Z}$ defined by
$\bar{b}(\bar{x},\bar{y})=b(x,y)\mod 1.$
Furthermore, when $L$ is even, the self-product $\bar{b}(\bar{x},\bar{x})$ is well-defined as an element of $\mathbb{Q}/2\mathbb{Z}$. Any isometry $t$ of $L$ induces an automorphism $\bar{t}:G(L)\to G(L)$ that preserves the torsion form $\bar{b}$.
Gluing
Let $L=(L,b)$ be a lattice, and $S$ a submodule of $L$. We say that $S$ is primitive in $L$ if $L/S$ is torsion-free, or equivalently, if $S=L\cap\mathbb{Q}S$ in $L\otimes\mathbb{Q}$. We define $S^{\perp}\coloneqq\{y\in L\mid b(y,x)=0\text{ for all $x\in S$}\}$. Note that such a submodule is always primitive.
Proposition 2.2.
Let $(L_{1},b_{1})$ and $(L_{2},b_{2})$ be two even lattices. Suppose that there exists an isomorphism $\gamma:G(L_{1})\to G(L_{2})$ such that
$\overline{b_{1}}(\bar{x},\bar{x})+\overline{b_{2}}(\gamma(\bar{x}),\gamma(\bar{x}))=0\quad\text{in $\mathbb{Q}/2\mathbb{Z}$ for all $\bar{x}\in G(L_{1})$}.$
Then there exists an even unimodular lattice $L$ such that $L_{1}$ and $L_{2}$ are primitive sublattices in $L$ with $L_{2}=L_{1}^{\perp}$.
Assume further that isometries $t_{1}$ and $t_{2}$ of $L_{1}$ and $L_{2}$ with $\gamma\circ\overline{t_{1}}=\overline{t_{2}}\circ\gamma$ are given. Then there exists an isometry $t$ of $L$ such that $t|_{L_{1}}=t_{1}$ and $t|_{L_{2}}=t_{2}$.
Proof.
See [9, page 5]. ∎
In Proposition 2.2, we remark that if $(r_{1},s_{1})$ and $(r_{2},s_{2})$ are the signatures of $L_{1}$ and $L_{2}$ respectively then that of $L$ is given by $(r_{1}+r_{2},s_{1}+s_{2})$; and if $F_{1}$ and $F_{2}$ are the characteristic polynomials of $t_{1}$ and $t_{2}$ respectively then that of $t$ is given by $F_{1}F_{2}$.
Twists
Let $L=(L,b)$ be a lattice, and $t:L\to L$ an isometry. We write $\mathbb{Z}[t+t^{-1}]$ for the subring of $\operatorname{End}(L)$ generated by $t+t^{-1}$. Let $a\in\mathbb{Z}[t+t^{-1}]$ be an element with $\det(a)\neq 0$. Then
$b_{a}(x,y)=b(ax,y)\quad(x,y\in L)$
defines a new inner product on $L$. The lattice $(L,b_{a})$ is called the twist of $L$ by $a$, and denoted by $L(a)$. We have $\det(L(a))=\det(a)\det(L)$. The isometry $t$ of $L$ is also an isometry of $L(a)$. If $L$ is even, then so is $L(a)$, see [9, Proposition 4.1].
Theorem 2.3.
Let $(L,b)$ be a lattice, $t$ an isometry of $L$, and $F\in\mathbb{Z}[X]$ the characteristic polynomial of $t$. Let $p$ be a prime not dividing $\det L$, and let $a\in\mathbb{Z}[t+t^{-1}]$ be an element with $\det(a)\neq 0$ and $pL\subset aL$. Suppose that $F\bmod p\in\mathbb{F}_{p}[X]$ is separable. Then
$G(L(a))\cong G(L)\oplus G(L(a))_{p}\quad\text{as $\mathbb{Z}[t]$-modules}$
and $p.G(L(a))_{p}=0$. Moreover, regarding $G(L(a))_{p}$ as an $\mathbb{F}_{p}$-vector space, the characteristic polynomial of $\bar{t}|_{G(L(a))_{p}}:G(L(a))_{p}\to G(L(a))_{p}$ is $\gcd(A(X)\bmod p,F(X)\bmod p)$, where $A(X)\in\mathbb{Z}[X]$ is a polynomial satisfying $a=A(t)$.
Proof.
See [9, Theorem 4.2]. ∎
3 Proof of the main theorems
We prove Theorems A and B. We refer to [7] for fundamental results on K3 surfaces.
3.1 Proof of Theorem A
Put $Q(X)=X^{2}-3X+1$. To prove Theorem A, we construct an isometry of a K3 lattice with characteristic polynomial $Q\cdot\Phi_{50}$ by gluing. Let us begin by constructing two lattice isometries whose characteristic polynomials are $Q$ and $\Phi_{50}$ respectively, ensuring that they are compatible for gluing. In the following, the prime number $3001$ appears. It is a factor of the resultant $\operatorname{Res}(Q,\Phi_{50})$, and is referred to as a feasible prime for $Q$ in [10].
Lemma 3.1.
Let $L=(\mathbb{Z}^{2},b)$ be the lattice of rank $2$ having Gram matrix $3001\begin{pmatrix}2&1\\ 1&-2\end{pmatrix}$ with respect to the standard basis. Then, it has signature $(1,1)$, and $G(L)\cong(\mathbb{Z}/5\mathbb{Z})\oplus(\mathbb{Z}/3001\mathbb{Z})^{2}$. Moreover $G(L)_{5}$ has a generator $\bar{v}$ such that $\bar{b}(\bar{v},\bar{v})=2/5$ in $\mathbb{Q}/2\mathbb{Z}$.
Proof.
The former assertion is straightforward. Let $e_{1},e_{2}$ denote the standard basis of $L$, and put $v=\frac{2}{5}e_{1}+\frac{1}{5}e_{2}\in L\otimes\mathbb{Q}$. Then $b(v,e_{1})=3001$ and $b(v,e_{2})=0$. These imply that $v\in L^{\vee}$. Furthermore, $\bar{v}$ is a generator of $G(L)_{5}\cong\mathbb{Z}/5\mathbb{Z}$ since it has order $5$. Because $b(v,v)=3001\cdot 2/5$, we have $\bar{b}(\bar{v},\bar{v})=2/5$ in $\mathbb{Q}/2\mathbb{Z}$. ∎
Proposition 3.2.
Let $L=(\mathbb{Z}^{2},b)$ be the lattice in Lemma 3.1. Then $t=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$ is an isometry of $L$ with characteristic polynomial $Q(X)$. Moreover
  1. (i)
    $\bar{t}|_{G(L)_{5}}=-\mathrm{id}_{G(L)_{5}}$; and
  2. (ii)
    when we consider $G(L)_{3001}$ as an $\mathbb{F}_{3001}$-vector space, the characteristic polynomial of $\bar{t}|_{G(L)_{3001}}$ is $Q(X)=(X+121)(X-124)\in\mathbb{F}_{3001}[X]$.
Proof.
We have
$\begin{pmatrix}1&1\\ 1&2\end{pmatrix}^{\top}\begin{pmatrix}3001\cdot 2&3001\cdot 1\\ 3001\cdot 1&3001\cdot(-2)\end{pmatrix}\begin{pmatrix}1&1\\ 1&2\end{pmatrix}=\begin{pmatrix}3001\cdot 2&3001\cdot 1\\ 3001\cdot 1&3001\cdot(-2)\end{pmatrix}$,
which shows that $t$ is an isometry of $L$. It is clear that the characteristic polynomial of $t$ is $Q(X)=X^{2}-3X+1$.
(i). Let $v=\frac{2}{5}e_{1}+\frac{1}{5}e_{2}$ as in Lemma 3.1. Then
$tv=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}\begin{pmatrix}2/5\\ 1/5\end{pmatrix}=\begin{pmatrix}3/5\\ 4/5\end{pmatrix}=\begin{pmatrix}1\\ 1\end{pmatrix}-v.$
This implies that $\bar{t}|_{G(L)_{5}}=-\mathrm{id}_{G(L)_{5}}$ since $\bar{v}$ generates $G(L)_{5}$.
(ii). Let $L^{\prime}$ be a lattice with Gram matrix $\begin{pmatrix}2&1\\ 1&-2\end{pmatrix}$. Note that $t$ is also an isometry of $L^{\prime}$ with characteristic polynomial $Q(X)$, and that $L=L^{\prime}(3001)$. By applying Theorem 2.3 to $L^{\prime}$ and $t$ with $p=a=3001$, we see that the characteristic polynomial of $\bar{t}|_{G(L)_{3001}}$ is $Q(X)=(X+121)(X-124)\in\mathbb{F}_{3001}[X]$. ∎
Let $\Psi_{50}(Y)\in\mathbb{Z}[Y]$ be the trace polynomial of $\Phi_{50}(X)$, that is, a unique polynomial of degree $10=\deg(\Phi_{50})/2$ such that $\Phi_{50}(X)=X^{10}\Psi_{50}(X+X^{-1})$. Put $\zeta\coloneqq\exp\left(\frac{2\pi\sqrt{-1}}{50}\right)$ and $K\coloneqq\mathbb{Q}(\zeta)\cong\mathbb{Q}[X]/(\Phi_{50})$. The field $K$ admits an involution $\iota:K\to K$ uniquely determined by $\iota(\zeta)=\zeta^{-1}$. We define an inner product $b$ on the $\mathbb{Q}$-vector space $K$ by
$b(x,y)=\operatorname{Tr}_{K/\mathbb{Q}}\left(\frac{x\,\iota(y)}{\Psi_{50}^{\prime}(\zeta+\zeta^{-1})}\right)\quad(x,y\in K),$
where $\Psi_{50}^{\prime}(Y)=\frac{d}{dY}\Psi_{50}(Y)$. Then, it follows from [8, Theorem 8.1] that $L\coloneqq(\mathcal{O}_{K},b)$ is an even lattice with $|\det L|=|\Phi_{50}(1)\Phi_{50}(-1)|=5$, where $\mathcal{O}_{K}$ is the ring of integers of $K$, which is equal to $\mathbb{Z}[\zeta]$. Furthermore, the linear transformation $t$ defined as the multiplication by $\zeta$ is an isometry of $L$ with characteristic polynomial $\Phi_{50}(X)$. Note that $\Phi_{50}(X)\bmod 3001$ is separable. In fact, for any positive integer $n$ and prime number $p$, $X^{n}-1\bmod p$ is separable if $p\nmid n$.
We will use a twist of this lattice $L$. For some computations below, we make use of computer algebra. We define $k\coloneqq\{x\in K\mid\iota(x)=x\}\subset K$. Then $k=\mathbb{Q}(\zeta+\zeta^{-1})$, and $\mathbb{Z}[t+t^{-1}]=\mathbb{Z}[\zeta+\zeta^{-1}]=\mathcal{O}_{k}$. Put
$u_{1}\coloneqq\zeta^{2}+1+\zeta^{-2},\quad u_{2}\coloneqq\sum_{i=0}^{5}(\zeta^{2i+1}+\zeta^{-(2i+1)}),\quad a^{\prime}\coloneqq\frac{\zeta+\zeta^{-1}-3}{\zeta+\zeta^{-1}+2},$
and
$a\coloneqq u_{1}u_{2}a^{\prime}=(\zeta^{2}+1+\zeta^{-2})\cdot\left(\sum_{i=0}^{5}(\zeta^{2i+1}+\zeta^{-(2i+1)})\right)\cdot\frac{\zeta+\zeta^{-1}-3}{\zeta+\zeta^{-1}+2}.$
One can check that $u_{1},u_{2}\in\mathcal{O}_{k}^{\times}$, and $a^{\prime}\in\mathcal{O}_{k}$. Hence $a\in\mathcal{O}_{k}$. Moreover, we have
$\operatorname{N}_{k/\mathbb{Q}}(a)=\operatorname{N}_{k/\mathbb{Q}}(a^{\prime})=\Psi_{50}(3)/\Psi_{50}(-2)=3001,$ (3)
where $\operatorname{N}_{k/\mathbb{Q}}$ is the norm map. This implies that $3001\mathcal{O}_{k}\subset a\mathcal{O}_{k}$, and thus $3001L\subset aL$. Therefore, we can apply Theorem 2.3 to the current $L$, $t$, and $a$, taking $p=3001$.
Proposition 3.3.
The twist $L(a)=(\mathcal{O}_{K},b_{a})$ of $L$ by $a$ is an even lattice of signature $(2,18)$. Moreover, the following assertions hold:
  1. (i)
    $G(L(a))\cong(\mathbb{Z}/5\mathbb{Z})\oplus(\mathbb{Z}/3001\mathbb{Z})^{2}$.
  2. (ii)
    $G(L(a))_{5}$ has a generator $\bar{v}$ such that $\overline{b_{a}}(\bar{v},\bar{v})=-2/5$ in $\mathbb{Q}/2\mathbb{Z}$.
  3. (iii)
    $\bar{t}|_{G(L(a))_{5}}=-\mathrm{id}_{G(L(a))_{5}}$.
  4. (iv)
    When we consider $G(L(a))_{3001}$ as an $\mathbb{F}_{3001}$-vector space, the characteristic polynomial of $\bar{t}|_{G(L(a))_{3001}}$ is $Q(X)=(X+121)(X-124)\in\mathbb{F}_{3001}[X]$.
Proof.
As mentioned in §2, the twist $L(a)$ is even since so is $L$. Furthermore, it follows from [5, Theorem 4.2] that the positive signature $r$ of $L(a)$ is given by
$r=2\cdot\#\{\sigma\in\operatorname{Gal}(k/\mathbb{Q})\mid\sigma(a/\Psi_{50}^{\prime}(\zeta+\zeta^{-1}))>0\}.$
Using computer algebra, one can check that $\sigma\in\operatorname{Gal}(k/\mathbb{Q})$ with $\sigma(\zeta+\zeta^{-1})=\zeta^{7}+\zeta^{-7}$ is the only element satisfying $\sigma(a/\Psi_{50}^{\prime}(\zeta+\zeta^{-1}))>0$, see Table 3.1. Hence $r=2$, and the signature of $L(a)$ is $(2,18)$.
$\sigma(\zeta+\zeta^{-1})$ $\zeta+\zeta^{-1}$ $\zeta^{3}+\zeta^{-3}$ $\zeta^{7}+\zeta^{-7}$ $\zeta^{9}+\zeta^{-9}$ $\zeta^{11}+\zeta^{-11}$
$\sigma(a/\Psi_{50}^{\prime}(\zeta+\zeta^{-1}))$ 0.11372-0.11372 0.067094-0.067094 $0.028027$ 0.026605-0.026605 0.11141-0.11141
$\sigma(\zeta+\zeta^{-1})$ $\zeta^{13}+\zeta^{-13}$ $\zeta^{17}+\zeta^{-17}$ $\zeta^{19}+\zeta^{-19}$ $\zeta^{21}+\zeta^{-21}$ $\zeta^{23}+\zeta^{-23}$
$\sigma(a/\Psi_{50}^{\prime}(\zeta+\zeta^{-1}))$ 0.10565-0.10565 0.029497-0.029497 0.5185-0.5185 1.5061-1.5061 2.5493-2.5493
Table 3.1: Approximate values of $\sigma(a/\Psi_{50}^{\prime}(\zeta+\zeta^{-1}))$. Note that $\sigma\in\operatorname{Gal}(k/\mathbb{Q})$ is determined by its value at $\zeta+\zeta^{-1}$.
(i). Since $|\det L|=5$, we have $G(L)=\mathbb{Z}/5\mathbb{Z}$. Moreover, we have $\operatorname{N}_{K/\mathbb{Q}}(a)=\operatorname{N}_{k/\mathbb{Q}}(a^{2})=3001^{2}$ by (3), and thus,
$|\det(L(a))|=|\det(a)\det(L)|=|\operatorname{N}_{K/\mathbb{Q}}(a)\det(L)|=3001^{2}\cdot 5.$
Hence, the assertion follows from the former statement of Theorem 2.3.
(ii). Put
$v=\frac{1}{5}(2+3\zeta+2\zeta^{2}+3\zeta^{3}+2\zeta^{4}+\zeta^{5}-\zeta^{6}+\zeta^{7}-\zeta^{8}+\zeta^{9}+\zeta^{10}\\ -\zeta^{11}+\zeta^{12}-\zeta^{13}+\zeta^{14}-3\zeta^{15}-2\zeta^{16}-3\zeta^{17}-2\zeta^{18}-3\zeta^{19}).$
Then, using computer algebra, one can verify that $b_{a}(v,\zeta^{j})$ is an integer for $j=0,\ldots,19$, and $b_{a}(v,v)=-142/5$. Hence, $v\in L(a)^{\vee}$, and $\bar{v}$ is the desired generator of $G(L(a))_{5}$.
(iii). We have $(t+\mathrm{id}_{L(a)^{\vee}})^{20}.L(a)^{\vee}\subset 5L(a)^{\vee}$ since $\Phi_{50}(X)\bmod 5=(X+1)^{20}$. Thus $(\bar{t}|_{G(L(a))_{5}}+\mathrm{id}_{G(L(a))_{5}})^{20}.G(L(a))_{5}=0$, which leads to
$(\bar{t}|_{G(L(a))_{5}}+\mathrm{id}_{G(L(a))_{5}}).G(L(a))_{5}=0$
since $G(L(a))_{5}$ is a one-dimensional $\mathbb{F}_{5}$-vector space. Hence $\bar{t}|_{G(L(a))_{5}}=-\mathrm{id}_{G(L(a))_{5}}$.
(iv). Note that $(X+121)(X-124)\mid(\Phi_{50}(X)\bmod 3001)$ in $\mathbb{F}_{3001}[X]$. Let $A(X)\in\mathbb{Z}[X]$ be a polynomial satisfying $a=A(\zeta)$. Since
$a^{\prime}=\frac{\zeta+\zeta^{-1}-3}{\zeta+\zeta^{-1}+2}=\frac{\zeta^{2}-3\zeta+1}{\zeta^{2}+2\zeta+1}=\frac{Q(\zeta)}{(\zeta+1)^{2}},$
we can write
$A(X)=U_{1}(X)U_{2}(X)h(X)Q(X)\mod\Phi_{50}(X),$
where $U_{i}(X)$ is a polynomial with $u_{i}=U_{i}(\zeta)$ ($i=1,2$) and $h(X)$ is the inverse of $(X+1)^{2}$ in $\mathbb{Z}[X]/(\Phi_{50})$. Hence $(Q\bmod 3001)$ is a common divisor of $(A\bmod 3001)$ and $(\Phi_{50}\bmod 3001)$ in $\mathbb{F}_{3001}[X]$. Since $G(L(a))_{3001}$ is two-dimensional over $\mathbb{F}_{3001}$, The latter assertion of Theorem 2.3 shows that the characteristic polynomial of $\bar{t}|_{G(L(a))_{3001}}$ is $Q(X)$. ∎
Remark 3.4.
One can show Proposition 3.3 by using matrix representation. For reference, if $G=(g_{ij})_{ij}$ is the Gram matrix of $L(a)$ with respect to the basis $1,\zeta,\zeta^{2},\ldots,\zeta^{19}$, then each entry $g_{ij}$ depends only on the difference iji-j, and the first row of $G$ is given by
(10,8,6,3,1,2,3,3,3,3,3,3,3,3,3,3,3,3,3,3).(-10,8,-6,3,-1,-2,3,-3,3,-3,3,-3,3,-3,3,-3,3,-3,3,-3).
We glue the two lattices in Propositions 3.2 and 3.3.
Theorem 3.5.
There exists an isometry $t$ of a K3 lattice $\Lambda$ with characteristic polynomial $Q\Phi_{50}$ such that $\Lambda$ contains a $t$-invariant primitive sublattice $L_{1}$ having Gram matrix $3001\begin{pmatrix}2&1\\ 1&-2\end{pmatrix}$ and the characteristic polynomial of $t|_{L_{1}}$ is $Q$.
Proof.
Let $(L_{1},b_{1})$ and $t_{1}$ (resp. $(L_{2},b_{2})$ and $t_{2}$) be the lattice and its isometry in Proposition 3.2 (resp. Proposition 3.3). There exists an isomorphism $\gamma_{5}:G(L_{1})_{5}\to G(L_{2})_{5}$ with $\overline{b_{1}}(\bar{x},\bar{x})+\overline{b_{2}}(\gamma_{5}(\bar{x}),\gamma_{5}(\bar{x}))=0$ in $\mathbb{Q}/2\mathbb{Z}$ for any $\bar{x}\in G(L_{1})_{5}$ by Lemma 3.1 and (ii) of Proposition 3.3. Furthermore $\gamma_{5}\circ\overline{t_{1}}|_{G(L_{1})_{5}}=-\gamma_{5}=\overline{t_{2}}|_{G(L_{2})_{5}}\circ\gamma_{5}$. On the other hand, it follows from [9, Theorem 3.1] together with (ii) of Proposition 3.2 and (iv) of Proposition 3.3 that there exists an isomorphism $\gamma_{3001}:G(L_{1})_{3001}\to G(L_{2})_{3001}$ such that $\overline{b_{1}}(\bar{x},\bar{x})+\overline{b_{2}}(\gamma_{3001}(\bar{x}),\gamma_{3001}(\bar{x}))=0$ in $\mathbb{Q}/2\mathbb{Z}$ for any $\bar{x}\in G(L_{1})_{3001}$ and $\gamma_{3001}\circ\overline{t_{1}}|_{G(L_{1})_{3001}}=\overline{t_{2}}|_{G(L_{2})_{3001}}\circ\gamma_{3001}$. Hence, the resulting isometry $t$ obtained by applying Proposition 2.2 to $\gamma_{5}\oplus\gamma_{3001}:G(L_{1})\to G(L_{2})$ is the desired one. The proof is complete. ∎
We are now ready to prove Theorem A.
Proof of Theorem A.
Let $t$ be the isometry of a K3 lattice $\Lambda$ in Theorem 3.5. Put $L_{2}=L_{1}^{\perp}$. Note that $L_{2}$ is also $t$-invariant, and the characteristic polynomial of $t|_{L_{2}}$ is $\Phi_{50}$. The signature of $L_{2}$ is $(2,18)$ since that of $L_{1}$ is $(1,1)$. This implies that there exists a root $\delta$ of $\Phi_{50}$ such that the signature of the subspace $\{v\in L\otimes\mathbb{R}\mid(t^{2}-(\delta+\delta^{-1})t+1).v=0\}$ is $(2,0)$. Let $\omega\in L\otimes\mathbb{C}$ be a nonzero eigenvector of $t$ corresponding to $\delta$. Then, by surjectivity of the period mapping (see [7, Theorem 6.9]), there exists a K3 surface $S$ and an isomorphism $\alpha:H^{2}(S,\mathbb{Z})\to\Lambda$ of lattices such that $\alpha(\omega_{S})=\omega$. We have
$\alpha(P_{S})=\alpha(\{x\in H^{2}(S,\mathbb{Z})\mid\langle x,\omega_{S}\rangle=0\})=L_{2}^{\perp}=L_{1},$
which implies that $\rho_{S}=\operatorname{rk}L_{1}=2$. It also implies that $P_{S}$ is indefinite, and thus, $S$ is projective (see [7, Proposition 4.11]). Moreover, $\alpha^{-1}\circ t\circ\alpha$ preserves the Kähler cone because $P_{S}\cong L_{1}$ has no element with self-intersection number 2-2. Hence, by the Torelli theorem (see [7, Theorem 6.1]), there exists an automorphism $\phi$ of $S$ such that $\phi^{*}=\alpha^{-1}\circ t\circ\alpha$. This completes the proof. ∎
3.2 Proof of Theorem B
Let $S$ be a projective K3 surface, and $\phi:S\to S$ an automorphism of $S$. We write $F(X)$ and $g(X)$ for the characteristic polynomials of $\phi^{*}:H^{2}(S)\to H^{2}(S)$ and $\phi^{*}|_{P_{S}}:P_{S}\to P_{S}$. It follows from [7, Corollary 8.13] that $F$ can be written as
$F(X)=g(X)\Phi_{l}(X)^{m}$ (4)
for some $l$ and $m\in\mathbb{Z}_{\geq 1}$, where $\Phi_{l}(X)$ is the $l$-th cyclotomic polynomial.
We now assume that $\rho_{S}=2$. Then, Equation (4) implies that $m\varphi(l)=22-\deg g=20$, where $\varphi(l):=\#(\mathbb{Z}/l\mathbb{Z})^{\times}=\deg\Phi_{l}$. Hence, $m$ and $\varphi(l)$ are divisors of $20$, and in particular,
$l\in\{1,2,3,4,6,5,8,10,12,11,22,25,33,44,50,66\}.$
We further assume that the dynamical degree $\lambda(\phi)$ is greater than $1$. Then $\lambda(\phi)$ is a Salem number of degree $2$, and we can write $g(X)=X^{2}-\tau X+1$, where $\tau\coloneqq\lambda(\phi)+\lambda(\phi)^{-1}\in\mathbb{Z}_{\geq 3}$.
Proposition 3.6.
We have $l=1,2,5,10,25,$ or $50$. Moreover, putting
$\epsilon=\begin{cases}1&\text{if $l=1,5,25$}\\ -1&\text{if $l=2,10,50$,}\end{cases}$ (5)
the integer $\tau+2\epsilon$ is a square number, and if $l\neq 1,2$ then $5(\tau-2\epsilon)$ is a square number.
Proof.
Since $H^{2}(S,\mathbb{Z})$ is an even unimodular lattice, it follows from Theorem 2.1 that
$|F(1)|=(\tau-2)|\Phi_{l}(1)|^{m}\quad\text{and}\quad|F(-1)|=(\tau+2)|\Phi_{l}(-1)|^{m}$
are squares. Suppose that $m=20$. Then $\varphi(l)=1$, which means that $l=1$ or $2$. If $l=1$ then $|F(-1)|=(\tau+2)\cdot|-2|^{20}$, and $\tau+2$ must be a square. If $l=2$ then $|F(1)|=(\tau-2)\cdot 2^{20}$, and $\tau-2$ must be a square. Suppose then that $m<20$. If $m$ were even then both $\tau+2$ and $\tau-2$ would be squares, but it is impossible. Hence $m$ is odd, which yields $m=1$ or $5$ since $m$ is a divisor of $20$.
Suppose that $m=5$. Then $\varphi(l)=4$, which means that $l\in\{5,8,10,12\}$. If $l$ were $8$ then $|F(1)|=(\tau-2)\Phi_{8}(1)^{5}=(\tau-2)\cdot 2^{5}$ and $|F(-1)|=(\tau+2)\Phi_{8}(-1)^{5}=(\tau+2)\cdot 2^{5}$, and therefore $2(\tau-2)$ and $2(\tau+2)$ would be squares. However, such a $\tau\in\mathbb{Z}_{\geq 3}$ does not exist. Hence $l\neq 8$. Similarly, we can get $l\neq 12$. If $l=5$ then $|F(1)|=(\tau-2)\Phi_{5}(1)^{5}=(\tau-2)\cdot 5^{5}$ and $|F(-1)|=(\tau+2)\Phi_{5}(-1)^{5}=\tau+2$, and hence $\tau+2$ and $5(\tau-2)$ are squares. Similarly, in the case $l=10$, we can show that $\tau-2$ and $5(\tau+2)$ are squares.
Suppose that $m=1$. Then $\varphi(l)=20$, which means that $l\in\{25,33,44,50,66\}$. By similar arguments as above, the cases $l=33$, $44$, and $66$ are ruled out, and it can be shown that: if $l=25$ then $\tau+2$ and $5(\tau-2)$ are squares; and if $l=50$ then $\tau-2$ and $5(\tau+2)$ are squares. The proof is complete. ∎
Let $\epsilon\in\{1,-1\}$ be as in Proposition 3.6, and let $\alpha\in\mathbb{Z}_{\geq 0}$ be the non-negative integer satisfying
$\tau+2\epsilon=\alpha^{2}.$
Since $\tau\geq 3$, if $\epsilon=1$ then $\alpha\geq 3$, and if $\epsilon=-1$ then $\alpha\geq 1$. This implies that
$\mathcal{T}\subset\{2\}\cup\{\gamma^{2}-2\mid\gamma\in\mathbb{Z}_{\geq 3}\}\cup\{\gamma^{2}+2\mid\gamma\in\mathbb{Z}_{\geq 1}\},$ (6)
where $\mathcal{T}$ is defined in (1). Note that the sets $\{\gamma^{2}-2\mid\gamma\in\mathbb{Z}_{\geq 3}\}$ and $\{\gamma^{2}+2\mid\gamma\in\mathbb{Z}_{\geq 1}\}$ are disjoint.
Lemma 3.7.
Suppose that $\epsilon=-1$. Then $\alpha\notin\{2,3,5,7,13,17\}$.
Proof.
We have $l=2$, $10$, or $50$ since $\epsilon=-1$. If $l=2$ then $\phi^{*}\omega_{S}=-\omega_{S}$, and thus $\alpha\notin\{2,3,5,7,13,17\}$ by Theorem 1.1. Suppose that $l=10$ or $50$. Then, $5(\tau+2)=5(\alpha^{2}+4)$ is a square number by Proposition 3.6. In particular, $\alpha^{2}+4$ is divisible by $5$. Hence, $\alpha$ cannot be any of $2,3,5,7,13,17$. ∎
Proof of Theorem B.
By (6) and Lemma 3.7, we have
$\mathcal{T}\subset\{2\}\cup\{\alpha^{2}-2\mid\alpha\in\mathbb{Z}_{\geq 3}\}\cup\{\alpha^{2}+2\mid\alpha\in\mathbb{Z}_{\geq 1}\setminus\{2,3,5,7,13,17\}\}.$
Together with the reverse inclusion (2), this completes the proof of Theorem B. ∎
Acknowledgments
This work is supported by JSPS KAKENHI Grant Number JP24KJ0044.